# A No-Go Theorem for Shaping Quantum Resources

### v2

@@ Abstract
We now formalize this impossibility as a theorem that identifies the symplectic subalgebra as the sole analytic structure closing the moment hierarchy.
@@
Corollary 1 (Rigidity of the Moment Hierarchy).
No independent control of higher cumulants) For any connected Lie flow generated by smooth Hamiltonians on phase space, the evolution of higher normalized moments is not independent of the first two. Equivalently, no Hamiltonian generator—finite or infinite dimensional—can alter a higher cumulant while leaving mean and covariance invariant for all states.
@@
Beyond its geometric interpretation, the rigidity theorem also provides a natural bridge between phase-space analysis and symplectic geometry. The second-order truncation condition corresponds to the vanishing of all third- and higher-order terms in the Moyal tensor hierarchy, so that the Gaussian manifold forms a flat symplectic submanifold endowed with a covariantly constant metric. Non-quadratic Hamiltonians introduce curvature proportional to the third derivatives of $H(x,p)$, giving a quantitative measure of departure from symplectic flatness. This curvature viewpoint connects analytic rigidity to the language of differential geometry and deformation quantization.
@@
Interpretation – Theorem 2 establishes a precise equivalence between three a priori different statements about smooth Hamiltonian dynamics: (i) the closure of the system of equations for the first and second statistical moments; (ii) the invariance of the finite polynomial space $\mathcal{P}_{\leq 2}$ under the adjoint generator $L_{H}^{\dagger}$; and (iii) the quadratic form of the Hamiltonian. This equivalence is not an approximation or semiclassical limit: it follows exactly from the structure of the Moyal series and holds for all $\hbar>0$. Formally, the hierarchy of normalized moments constitutes an infinite tower of coupled ordinary differential equations in expectation values. Theorem 2 states that the only smooth Hamiltonian vector fields on phase space that leave the lower two tiers of this tower self-contained are those generated by quadratic forms. Every smooth non-quadratic term creates a non-vanishing component of $L_{H}^{\dagger}f$ orthogonal to $\mathcal{P}_{\leq 2}$, thereby enforcing coupling to higher cumulants. Physically, this means that no Hamiltonian evolution can modify higher-order statistics without simultaneously perturbing the mean and covariance: the Gaussian manifold is an analytically rigid submanifold of the full state space. The symplectic algebra $\mathfrak{sp}(2N,\mathbb{R})$ thus plays the same role in continuous-variable mechanics that the Clifford algebra plays in discrete quantum computation: it is the maximal finite-order structure whose action preserves the algebraic closure of observables and admits efficient classical simulation.
@@
Broader Implications – Beyond verification, the theorem recasts several pillars of continuous-variable quantum information. In resource-theoretic language, the Gaussian manifold defines the set of free operations, and rigidity forbids leaving this set without simultaneously altering the Gaussian statistics. Monotones based on skewness, kurtosis, or higher cumulants are therefore not independent; they evolve coherently along low-dimensional trajectories dictated by the underlying symplectic symmetry. For bosonic error-correcting codes this means that Hamiltonian-only encodings cannot reshape the distribution tails independently of the bulk. Cat, binomial, and GKP codes [8, 9, 10, 11] all rely on higher-moment shaping that, by the theorem, requires measurement or ancilla assistance. In continuous-variable QKD [12, 13], variance and kurtosis often appear as separable noise parameters, but rigidity removes that freedom: no Hamiltonian process can vary one without the other, simplifying security analyses and eliminating hypothetical attacks based on unattainable independent cumulant control.
@@
The theorem invites several extensions. Can similar rigidity principles appear in discrete or hybrid systems governed by smooth group actions on probability simplices? Is there a resource-theoretic metric that measures the curvature of the Gaussian manifold directly from experimental data? And can engineered measurement-based feedback circumvent analytic rigidity by introducing non-smooth dynamics? Pursuing these questions will clarify how the analytic structure of quantum mechanics delineates not only what can be simulated but also what can be physically realized.

### v3

@@ Abstract
We now formalize this impossibility as a theorem that identifies the symplectic subalgebra as the sole analytic structure closing the moment hierarchy. Throughout, by a smooth Hamiltonian we mean a phase-space function $H(q,p)$ belonging to a differentiability class sufficient to ensure that the associated Moyal generator admits a finite-order differential representation. In particular, nonquadratic contributions require differentiability to at least third order, and for definiteness we assume $H\in C^{\infty}$.
@@
We emphasize that the theorem concerns dynamical invariance of the first and second moments under Hamiltonian evolution. While any non-Gaussian state can be associated with a Gaussian state sharing the same mean and covariance, such post hoc matching does not imply that a nonquadratic Hamiltonian preserves these moments along its dynamical flow. For cubic Hamiltonians (e.g., $H\propto q^{3}$), the time derivatives of the first and second moments are generically nonzero, reflecting the structural coupling of the moment hierarchy established above.
@@
Corollary 1 (No independent control of higher cumulants).
For any connected Lie flow generated by smooth Hamiltonians on phase space, the evolution of higher normalized moments is not independent of the first two. Equivalently, no Hamiltonian generator—finite or infinite dimensional—can alter a higher cumulant while leaving mean and covariance invariant for all states.
@@
Beyond its geometric interpretation, the rigidity theorem also provides a natural bridge between phase-space analysis and symplectic geometry – the second-order truncation condition corresponds to the vanishing of all third and higher-order terms in the Moyal tensor hierarchy, so that the Gaussian manifold forms a flat symplectic submanifold endowed with a covariantly constant metric. Non-quadratic Hamiltonians introduce curvature proportional to the third derivatives of $H(x,p)$, giving a quantitative measure of departure from symplectic flatness. This curvature viewpoint connects analytic rigidity to the language of differential geometry and deformation quantization.
@@
Interpretation – Theorem 2 establishes a precise equivalence between three seemingly different statements about smooth Hamiltonian dynamics: (i) the closure of the system of equations for the first and second statistical moments; (ii) the invariance of the finite polynomial space $\mathcal{P}_{\leq 2}$ under the adjoint generator $L_{H}^{\dagger}$; and (iii) the quadratic form of the Hamiltonian. This equivalence is not an approximation or semiclassical limit: it follows exactly from the structure of the Moyal series and holds for all $\hbar>0$. Formally, the hierarchy of normalized moments constitutes an infinite tower of coupled ordinary differential equations in expectation values. Theorem 2 states that the only smooth Hamiltonian vector fields on phase space that leave the lower two tiers of this tower self-contained are those generated by quadratic forms. Every smooth non-quadratic term creates a non-vanishing component of $L_{H}^{\dagger}f$ orthogonal to $\mathcal{P}_{\leq 2}$, thereby enforcing coupling to higher cumulants. Physically, this means that no Hamiltonian evolution can modify higher-order statistics without simultaneously perturbing the mean and covariance: the Gaussian manifold is an analytically rigid submanifold of the full state space. The symplectic algebra $\mathfrak{sp}(2N,\mathbb{R})$ thus plays the same role in continuous-variable mechanics that the Clifford algebra plays in discrete quantum computation: it is the maximal finite-order structure whose action preserves the algebraic closure of observables and admits efficient classical simulation.
The obstruction identified here is structural rather than parametric. Gaussian dynamics forms the unique nontrivial subspace with a closed moment hierarchy under Hamiltonian evolution. The inclusion of any higher-degree term—regardless of the number of independent nonlinear parameters—introduces higher-order differential contributions to the generator, thereby coupling lower and higher moments. Independent tuning of higher-order statistics cannot therefore be achieved merely by enlarging the parameter set.
@@
Broader Implications – This work recasts several pillars of continuous-variable quantum information. In resource-theoretic language, the Gaussian manifold defines the set of free operations, and rigidity forbids leaving this set without simultaneously altering the Gaussian statistics. Monotones based on skewness, kurtosis, or higher cumulants are therefore not independent; they evolve coherently along low-dimensional trajectories dictated by the underlying symplectic symmetry. For bosonic error-correcting codes this means that Hamiltonian-only encodings cannot reshape the distribution tails independently of the bulk. Cat, binomial, and GKP codes [8, 9, 10, 11] all rely on higher-moment shaping that, by the theorem, requires measurement or ancilla assistance. In continuous-variable QKD [12, 13], variance and kurtosis often appear as separable noise parameters, but rigidity removes that freedom: no Hamiltonian process can vary one without the other, simplifying security analyses and eliminating hypothetical attacks based on unattainable independent cumulant control.
@@
Our results also naturally raise several several questions. Can similar rigidity principles appear in discrete or hybrid systems governed by smooth group actions on probability simplices? Is there a resource-theoretic metric that measures the curvature of the Gaussian manifold directly from experimental data? And can engineered measurement-based feedback circumvent analytic rigidity by introducing non-smooth dynamics? Pursuing these questions will clarify how the analytic structure of quantum mechanics delineates not only what can be simulated but also what can be physically realized.